OpenStudy (anonymous):

Jack can paint a room in 6 ½ hours and Jill can paint a room in 4 ¾ hours. Jack and Jill can paint the room in A hours and B minutes if they work together. What is the value of B rounded to the nearest whole number? please show all your work.

OpenStudy (saifoo.khan):

Convert that into minutes first, then solve.

hero (hero):

x/6.5 + x/4.75 = 1

hero (hero):

Solve it first. You can convert later

OpenStudy (anonymous):

i hate mixed numbers. put 6 and a half hours as 390 minutes, 4 and a half is 285 minutes then your answer is \[\frac{390\times 285}{390+285}\] a calculator is next. then turn in to hours and minutes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but how can i get the value of B

hero (hero):

I know B = 47 minutes

hero (hero):

Good luck finding the steps

OpenStudy (anonymous):

the answer is 45

hero (hero):

What?

OpenStudy (anonymous):

yes the answer is 45

OpenStudy (anonymous):

It's from 8th grade math tournament.I have the answer.It's not from my calculator

hero (hero):

Okay, my calculator is broken